# Characterizing and Efficiently Accelerating Multimodal Generation Model Inference

### v1

@@ IV-A1 Accelerating Attention
To accelerate Attention, we enable PyTorch SDPA (Scaled Dot Product Attention) [29] that is designed specifically to accelerate the execution time of the fundamental building block — Attention — in Transformer-based model architectures. PyTorch provides torch.nn.functional.scaled_dot_product_attention as a function to optimize the inference time performance by accelerating the dot product computation between the Query, Key, and Value matrices using SDPA [46].
Instead of relying on the PyTorch SDPA API directly, for HSTU, we manually implemented the memory-efficient attention [32] and Flash Attention [10] as is in PyTorch, directly at HSTU’s internal code base. The memory-efficient attention implementation divides the input into blocks and avoid materializing the large $h\times N\times N$ intermediate attention tensors for the backward pass. This reduces the attention computation as a group of back-to-back GEMMs with different shapes, which enables the sparsity of input sequences to be exploited. The construction of the relative attention bias is also a bottleneck due to memory accesses. To address this issue, we fused the relative bias construction and grouped GEMMs into a single GPU kernel, and accumulates gradients using GPU’s fast shared memory in the backward pass.
Results – SPDA. Figure 6 presents the end-to-end inference latency speedup across the family of multi-modal generation tasks for the settings of batch size=1 and of the maximum batch size (the largest batch size that can support each model on a single A100 NVIDIA GPU as configured in Table III). PyTorch SDPA accelerates inference time performance of the generation tasks by an average of 1.07$\times$×and 1.43$\times$×for the single-batch and maximum-batch settings, respectively. In particular for HSTU, using the same fundamental principle, we observe 2.11$\times$ and 9.87$\times$ inference time improvement for the single-batch and maximum-batch settings, respectively. The significant inference time speedup stems from the proportionally-larger amount of time spent on the Attention operation for HSTU than the other generation tasks. And we observed that HSTU optimized implementation achieves up to 15$\times$ speedups on 8K sequences.
@@
Figure 5: End-to-end inference time speedup with SDPA and SDPA+torch.compile for Llama and Chameleon
@@
Figure 6: End-to-end inference time speedup with SDPA and SDPA+torch.compile for Seamless and HSTU, excluding AutoQuant optimization as linear operation is not a bottleneck according to Figure 4. Refer to Section IV-B for more details.
IV-A2 Improving GPU Utilization
During inference, especially for the single batch setting, the workloads are typically not compute bound, which raises 2 issues. First, each kernel that runs on the GPU becomes so fast, the overhead of launching kernels starts dominating the overall inference time. We reduce the number of kernels with PyTorch’s compiler. torch.compile [2] accelerates PyTorch models by capturing and optimizing their entire computation graph. This includes fusing multiple operations into a single kernel. The second and more important issue of inference is that the GPU computations can be faster than the time it takes to execute the corresponding python code on CPU. The consequence is that the GPU is inactive most of the time, waiting for instructions from the CPU. We address this with CUDA graphs [26]. A CUDA graph is a succession of GPU operations that can be executed as a whole, without having to execute CPU code to schedule kernels one-by-one. In particular, this ensures that the GPU is always active during the graph execution. In practice, the graph is captured once when running the PyTorch model, and can then be replayed whenever we have a new input.
@@ IV-A2 Improving GPU Utilization
Results – torch.compile/CUDA Graph. Figure 6 presents the additional inference performance speedup with torch.compile and CUDA Graph for the two settings. Overall, the end-to-end inference performance sees an average of 2.14$\times$ and 2.16$\times$ additional speedup on top of sdpa for the two batch settings, respectively. This results in total 2.28$\times$ and 3.09$\times$ speedup over the baseline without any optimization.
@@
Figure 7: End-to-end inference speedup of applying torch.compile and CUDA graph incrementally.
@@
A Deeper Dive with Seamless: Seamless is an emerging speech translation technology that is important to many product surfaces but has not received similar amount of attention as LLMs nor deep learning recommendation models. We focus significant performance acceleration efforts to enable real-time speech translation built on Seamless and present our key findings in a deeper dive here.
@@
As Figure 2(b) shows , there are four primary modules in Seamless. The T2TT decoder and vocoder are the most time-consuming modules, accounting for 61% and 23% of the end-to-end model inference time, respectively. Enabling torch.compile (mode=”max-autotune”) and CUDA Graph for the T2TT decoder and vocoder achieves 2$\times$ speedup for the text decoder and 30$\times$ speedup for the vocoder. This leads to 2.65$\times$ faster end-to-end inference latency. It turns out, in particular for single batch inference, GPU kernel launch time is hardly amortized, leading to substantial GPU idle time. Enabling torch.compile without CUDA Graph leaves the performance acceleration potential to to 1.17$\times$ and 18.4$\times$ for the text decoder and the vocoder, respectively. While still significant, it shows the important role of CUDA Graph.
@@
While torch.compile and CUDA Graph are key to inference time improvement, our detailed operator time breakdown in Figure 4 illustrates that Seamless spends significant amount of time on KV cache management (KV_Cache_Reorder). This is because Seamless adopts beam search as a text decoding strategy. In each incremental decoding step, beam search picks the ’N’ beams containing the best sequences so far based on the probabilities of the combination of all of the preceding words + current word. For each incremental decoding step, KV cache reordering is needed by all Attention layers to ensure that newly selected beams perform on the corresponding KV caches from previous decoding step — kv_cache = kv_cache.index_select(new_beams). This code allocates new memory space and overwrites the original memory pointer for kv_cache. In order to enable torch.compile for KV_Cache_Reorder, we had to modify KV cache reordering to keep the memory pointer of each cache as was recorded by using torch.Tensor.copy_ operator. By enabling torch.compile, all GPU kernelrs related to this operations are fused into a single function and compiled, resulting in the final speedup for Seamless.
@@
Figure 7 presents the overall inference speedup we achieve for Seamless step-by-step and Table V describes the each label used for the figure. While application-specific performance optimization, such as incremental decoding and KV cache reordering, is important, significant inference acceleration potential can be further achieved by torch.compile and CUDA Graph optimization. For Seamless M4T, an end-to-end inference speedup of 2.7$\times$ can be achieved for the challenging single-batch setting. This is key to efficiently enable low-latency, real-time speech translation tasks.
@@
[Text Decoder] Compile Apply torch.compile to the text decoder
[Text Decoder] Compile + CUDA Graph Apply torch.compile+CUDA Graph to the text decoder on top of the above row
+[KV Cache Reorder] Compile Apply torch.compile to KV cache reordering operation on top of the above row
@@
+[Vocoder] Compile + CUDA Graph Apply torch.compile + CUDA Graph to the vocoder on top of the above row
TABLE V: Table for description of the labels used in Figure 7.
IV-B Data Type Optimization
Quantization is an important optimization step before models are deployed for downstream inference. To understand the potential of state-of-the-art quantization capabilities and the impact on the baseline, we assess data type optimization by applying AutoQuant (Auto-Quantization)[30]. AutoQuant is a recently developed quantization implementation within the PyTorch torchao library [30] designed to integrate high-performance custom data types, layouts, and kernels into PyTorch workflows. AutoQuant optimizes the quantization process by determining the most efficient quantization method for each model layer. It supports two quantization types — int8 dynamic quantization, int8 weight-only quantization.
@@ IV-B Data Type Optimization
Results – AutoQuant. Figure 6 presents the inference time speedup when AutoQuant is enabled. Compared with the speedup with torch.compile (Section IV-A2), AutoQuant provides additional 1.20$\times$, 1.57$\times$ performance improvements for single batch setting. Compared to the baseline without any optimization, we observe an average of 2.13$\times$ and 4.38$\times$ latency improvement for the single and the maximum batch settings, respectively.
@@
For the other generation tasks using the model architectures of Seamless and HSTU, we do not expect performance improvement based on the operator time characterization results in Figure 4 — linear operations do not contribute significant runtime to end-to-end model inference. Furthermore, quantization optimization needs careful tuning, especially for production use cases of recommendation models [11]. Thus, we do not enable AutoQuant for HSTU.
IV-C Algorithm and Neural Network Specific Optimizations
@@
To meet the low inference latency requirement with resource efficiency in mind, we prioritize enabling system optimization levers that come with minimal accuracy impact — SDPA and Flash Attention in Section IV-A1 [16], torch.compile and CUDA Graph in Section IV-A2, and AutoQuant in Section IV-B. To further efficiently accelerate inference, algorithm and neural network specific optimizations are introduced. In this section, we focus on a state-of-the-art inference optimization technique: LayerSkip [12]. This technique is originally proposed to accelerate inference time for Llama. Here, we show how this techniques could be utilized to accelerate other multi-modal generative models.
@@
Figure 8: End-to-end inference time speedup with LayerSkip with batch size = 1.
LayerSkip [12] is specialized to minimize single-batch inference latency of LLMs for real-time, interactive conversations. During training, layer dropout [40] is applied, using low dropout rates for earlier layers and higher dropout rates for later layers with an early exit loss where all layers share the same classification head. Then, during inference, the training recipe increases the accuracy of early exit at the earlier layers, without adding any auxiliary layers or modules to the model. Finally, self-speculative decoding is introduced, where a subset of earlier layers are used to generate tokens sequentially, and remaining layers are used to verify and corrects tokens in parallel, amortizing the cost of loading their weights.
@@
Results – LayerSkip. To show the performance gain from workload-specific optimizations, LayerSkip, we choose Llama and Chameleon as our target models and show the inference time speedup in Figure 8. We focus on batch size 1 because efficient speculative decoding for larger batch sizes require significant modification to the attention mechanism [31, 8]. However, LayerSkip is an optimization technique that achieves significant inference time speedup at the cost of accuracy loss. We achieve 1.59$\times$ and 1.53$\times$ speedup with +2.5% and -1.2% accuracy impact for CodeLlama 7B and 34B model, respectively. For Chameleon 7B model, LayerSkip achieves 1.43$\times$ and 1.83$\times$ speedup with -3.2 and -6.36 cider score loss for I-T and IT-T tasks, respectively. Overall, we observed the geomean 1.58$\times$×speedup only with LayerSkip.
@@
Results – Putting It Altogether. We further explored performance gains by enabling all cross-stack optimization techniques, including system-level optimization techniques (SDPA, torch.compile, AutoQuant) and workload-load specific optimization technique (LayerSkip). This enhanced the speedup from 1.58$\times$ to 3.88$\times$, demonstrating the significant potential of combining these techniques for optimal performance gains.
V Key Lessons and Concluding Remarks

### v2

@@ 4.1.1 Accelerating Attention
To accelerate Attention, we enable PyTorch SDPA (Scaled Dot Product Attention) PyTorch (2023) designed specifically to accelerate the fundamental building block — Attention — in Transformer-based model architectures. PyTorch provides torch.nn.functional.scaled_dot_product_attention as a function to optimize the inference time performance by accelerating the dot product computation between the Query, Key, and Value matrices using SDPA Vaswani et al. (2023).
Instead of relying on the PyTorch SDPA API directly, for HSTU, we manually implemented the memory-efficient attention Rabe and Staats (2022) and Flash Attention Dao et al. (2022) as is in PyTorch, directly at HSTU’s internal code base. The memory-efficient attention implementation divides the input into blocks and avoid materializing the large $h\times N\times N$ intermediate attention tensors for the backward pass. This reduces the attention computation as a group of back-to-back GEMMs with different shapes, which enables the sparsity of input sequences to be exploited. The construction of the relative attention bias is also a bottleneck due to memory accesses. To address this issue, we fused the relative bias construction and grouped GEMMs into a single GPU kernel, and accumulates gradients using GPU’s fast shared memory in the backward pass.
Figure 5: End-to-end inference time speedup with SDPA and torch.compile for Llama and Chameleon on A100 GPU.
Results – SPDA. Figure 5 presents the end-to-end latency speedup across the family of multi-modal generation tasks for the settings of batch size=1 and of the maximum batch size (the largest batch size that can support each model on a single A100 NVIDIA GPU as configured in Table 3). PyTorch SDPA accelerates inference time performance of the generation tasks by an average of 1.07$\times$×and 1.43$\times$×for the single-batch and maximum-batch settings, respectively. In particular for HSTU, using the same fundamental principle, we observe 2.11$\times$ and 9.87$\times$ inference time improvement for the single-batch and maximum-batch settings, respectively. The significant inference time speedup stems from the proportionally-larger amount of time spent on the Attention operation for HSTU than the other generation tasks. And we observed that HSTU optimized implementation achieves up to 15$\times$ speedups on 8K sequences.
@@
Figure 6: End-to-end inference time speedup with SDPA and torch.compile for Seamless and HSTU on NVIDIA A100 GPU.
4.1.2 Improving GPU Utilization
During inference, for the single batch setting, the workloads are typically not compute bound, which raises 2 issues. First, each kernel that runs on the GPU becomes so fast, the overhead of launching kernels starts dominating the overall inference time. We reduce the number of kernels with PyTorch’s compiler. torch.compile Ansel et al. (2024) accelerates PyTorch models by capturing and optimizing their computation graph. This includes fusing multiple operations into a single kernel. The second and more important issue of inference is that the GPU computations can be faster than the time it takes to execute the corresponding python code on CPU. The consequence is that the GPU is inactive most of the time, waiting for instructions from the CPU. We address this with CUDA graphs NVIDIA (2018). A CUDA graph is a succession of GPU operations that can be executed as a whole, without having to execute CPU code to schedule kernels one-by-one. In particular, this ensures that the GPU is always active during the graph execution. In practice, the graph is captured once when running the PyTorch model, and can then be replayed whenever we have a new input.
@@ 4.1.2 Improving GPU Utilization
Results – torch.compile/CUDA Graph. Figure 6 presents the additional inference performance speedup with torch.compile and CUDA Graph. Overall, the end-to-end inference performance sees an average of 2.14$\times$ and 2.16$\times$ additional speedup on top of sdpa for the two batch settings, respectively. This results in total 2.28$\times$ and 3.09$\times$ speedup over the baseline without any optimization.
An exception, we observed performance degradation for Seamless at maximum batch size due to CUDA graph’s requirement for static KV cache. As mentioned earlier, substantial speedups are typically expected when static KV cache is used with CUDA Graph despite increased computational demands. However, this is an example of where the computational cost increase from static KV cache surpasses the benefits provided by CUDA Graph.
@@
Figure 7: End-to-end inference speedup of applying torch.compile incrementally on NVIDIA A100 GPU.
@@
A Deeper Dive with Seamless: Seamless is an emerging speech translation technology that is important to many product surfaces but has not received similar amount of attention as LLMs nor deep learning recommendation models. We focus significant performance acceleration efforts to enable real-time speech translation built on Seamless and present our key findings in a deeper dive.
@@
There are four primary modules in Seamless (Figure 2(b)). Enabling torch.compile (mode="max-autotune") and CUDA Graph for the T2TT decoder and vocoder (the most time-consuming modules, 61% and 23% of inference time) achieves 2$\times$ speedup for the text decoder and 30$\times$ speedup for the vocoder. This leads to 2.65$\times$ faster end-to-end inference latency. It turns out, in particular for single batch inference, GPU kernel launch time is hardly amortized, leading to substantial GPU idle time. Enabling torch.compile without CUDA Graph leaves the performance acceleration potential to to 1.17$\times$ and 18.4$\times$ for the text decoder and the vocoder, respectively. While still significant, it shows the important role of CUDA Graph.
@@
Our operator time breakdown in Figure 4 illustrates that Seamless also spends significant amount of time on KV cache management (KV_Cache_Reorder). This is because Seamless adopts beam search as a text decoding strategy. In each incremental decoding step, beam search picks the ’N’ beams containing the best sequences so far based on the probabilities of the combination of all of the preceding words + current word. For each incremental decoding step, KV cache reordering is needed by all Attention layers to ensure that newly selected beams perform on the corresponding KV caches from previous decoding step — kv_cache = kv_cache.index_select(new_beams). This code allocates new memory space and overwrites the memory pointer for kv_cache. To enable torch.compile for KV_Cache_Reorder, we had to modify KV cache reordering to keep the memory pointer of each cache as was recorded by using torch.Tensor.copy_ operator. By enabling torch.compile, all GPU kernels related to reordering are fused and compiled, resulting in the final speedup for Seamless.
@@
Figure 7 presents the overall inference speedup we achieve for Seamless step-by-step and Table 4 describes the each label used for the figure. While application-specific performance optimization, such as incremental decoding and KV cache reordering, is important, significant inference acceleration potential can be further achieved by torch.compile and CUDA Graph optimization. For Seamless M4T, an end-to-end inference speedup of 2.7$\times$ can be achieved for the challenging single-batch setting. This is key to efficiently enable low-latency, real-time speech translation tasks.
@@
[Text Dec. Compile Apply torch.compile to the text decoder
[Text Dec.] Compile + CUDA Graph Apply torch.compile+CUDA Graph to the text decoder on top of above row
+[KV Cache Reorder] Compile Apply torch.compile to KV cache reordering on top of above row
@@
+[Vocoder] Compile + CUDA Graph Apply torch.compile + CUDA Graph to the vocoder on top of above row
Table 4: Table for description of the labels used in Figure 7.
4.2 Data Type Optimization
Quantization is an important optimization before models are deployed for downstream inference. To understand the potential of quantization capabilities, we assess data type optimization by applying AutoQuant (Auto-Quantization)PyTorch (2024). AutoQuant is a recently developed quantization implementation within the PyTorch torchao library PyTorch (2024) designed to integrate high-performance custom data types, layouts, and kernels into PyTorch workflows. AutoQuant optimizes the quantization process by determining the most efficient quantization for each model layer. It supports two quantization types — int8 dynamic quantization, int8 weight-only quantization.
@@ 4.2 Data Type Optimization
AutoQuant is designed to work in conjunction with torch.compile, utilizing max-autotune setting to optimize quantization and achieve maximum performance gains. The quantization kernels within AutoQuant rely on torch.compile to generate high-performance kernels, therefore models must first be adapted to use static KV cache and static memory as highlighted in Section 4.1.2.
@@
Results – AutoQuant. Figure 6 presents the inference time speedup for AutoQuant. AutoQuant provides additional 1.20$\times$, 1.57$\times$ performance improvements for single batch setting on top of torch.compile (Section 4.1.2). Compared to the baseline without any optimization, we observe an average of 2.13$\times$ and 4.38$\times$ latency improvement for the single and the maximum batch settings, respectively.
@@
For other generation tasks using the model architectures of Seamless and HSTU, we do not expect performance improvement based on the characterization results in Figure 4 — linear operations do not contribute significant runtime to end-to-end model inference. Furthermore, quantization optimization needs careful tuning, especially for production use cases of recommendation models Deng et al. (2021), thus we opt out HSTU from AutoQuant enablement.
4.3 Algorithm and NN Specific Optimizations
@@
To meet the low inference latency requirement with resource efficiency, we prioritize enabling system optimization levers that come with minimal accuracy impact — SDPA and Flash Attention in Section 4.1.1 Golden et al. (2024b), torch.compile and CUDA Graph in Section 4.1.2, and AutoQuant in Section 4.2. To further efficiently accelerate inference, algorithm and neural network specific optimizations levers could be exploited. Here, we focus on a state-of-the-art inference optimization technique: LayerSkip Elhoushi et al. (2024). This technique is originally designed for Llama inference time and we show how it could be utilized to accelerate other multi-modal generative models.
@@
Figure 8: End-to-end inference time speedup with LayerSkip with batch size = 1 on NVIDIA A100 GPU.
LayerSkip Elhoushi et al. (2024) is specialized to minimize single-batch inference latency of LLMs. It speeds up inference by generating draft tokens sequentially with fewer layers while verifying them in parallel with remaining layers. Like speculative decoding Leviathan et al. (2023), parallel token verification amortizes per-token layer weight loading costs, resulting in end-to-end speedup. However, accuracy loss with exiting early is recovered by finetuning the model with a training recipe Elhoushi et al. (2024).
However, LayerSkip requires continuous model pretraining with early exit loss and layer dropout to improve early layer accuracy. It required 64 GPUs for 50K iterations for Code Llama and Chameleon, and moreover required access to the same or similar pretraining corpus that the models were trained on. Also, LayerSkip can be beneficial only for auto-regressive decoder models such as Llama, Chameleon, but not for non auto-regressive models like HSTU. Moreover, LayerSkip requires speculative decoding implementation for its mechanism, thus a custom implementation is required.
@@
Results – LayerSkip. Figure 8 shows the inference time performance gain from workload-specific optimizations, we choose Llama and Chameleon as our target models. We focus on batch size 1 because efficient speculative decoding for larger batch sizes require significant modification to the attention mechanism Qian et al. (2024); Daniel (2024). However, LayerSkip is an optimization technique that achieves significant inference time speedup at the cost of accuracy loss. We achieve 1.59$\times$ and 1.53$\times$ speedup with +2.5% and -1.2% accuracy impact for CodeLlama 7B and 34B model, respectively. For Chameleon 7B model, LayerSkip achieves 1.43$\times$ and 1.83$\times$ speedup with -3.2 and -6.36 cider score loss for I-T and IT-T tasks, respectively. Overall, we observed the geomean 1.58$\times$×speedup only with LayerSkip.
@@
Results – Putting It Altogether. We further explored performance gains by enabling all cross-stack optimization techniques, system-level optimizations (SDPA, torch.compile, AutoQuant) and workload-load specific optimization (LayerSkip). This enhanced the speedup from 1.58$\times$ to 3.88$\times$, demonstrating the significant potential of combining techniques for optimal performance gains.
4.4 Roofline Analysis
Figure 9 illustrates the effects of various optimization techniques on performance, as evaluated through the roofline analysis (data collected from NSight Compute NVIDIA (2024) profiling tool from NVIDIA). For each workload, Baseline is indicated with a circle marker where none of the optimization techniques is applied whereas Sys-Opt is indicated with a star marker where all the optimization levers are enabled. For Llama and Chameleon, SDPA+torch.compile+AutoQuant are enabled while SDPA+torch.compile is enabled for Seamless and SDPA is enabled for HSTU.
For each workload, enabling the system-level optimization techniques increased arithmetic intensity (i.e., FLOP/memory_traffic) and the performance (i.e., FLOP per sec), moving workload characteristics to the upper right part of the roofline. In the A100 deployment case, workloads that were already memory bandwidth-bound in the baseline setup are able to reduce the memory traffic and improve overall system performance. SDPA minimizes memory accesses during the attention mechanism by breaking down input sequence lengths into smaller tiles and perform computation within each tile independently. torch.compile fuses operations, eliminating intermediate memory allocations and accesses. AutoQuant effectively decreases the memory usage/traffic of each weight parameter by lowering their numerical precision.
The effect of applying each of these optimizations also depends on properties of each workload such as the underlying model architecture and input sequence length. For instance, we observe that workloads with textual inputs (e.g., T-T, T-I) and were previously most memory bandwidth-bound were the biggest beneficiaries. However, Seamless shows at most 10% difference (5% on average) between circle and star marker across four workloads (S-S, S-T, T-S, T-T). As mentioned in Section 4.1.2, applying torch.compile only to the text decoder among four primary modules in Seamless results in trivial impact to overall arithmetic intensity and the performance.
Figure 9: Roofline Analysis for Generative AI Workloads.
Beyond the Roofline Analysis We take Llama as an example to further understand the performance implication of each optimization technique to the roofline. When we apply SDPA, two key effects occur: First, FLOPs count increases by 8%. This is because efficient attention techniques require some recomputation. Second, memory traffic decreases by 14% due to the optimized algorithm. As a result, the arithmetic intensity increases. Counterintuitively, applying torch.compile on top of SDPA both increases both FLOPs count and memory traffic due to static KV cache adoption. Overall arithmetic intensity still increases because FLOPs count increases at a faster rate compared to memory traffic (attention has $O(N^{2})$ complexity). Memory traffic increases slightly by 1%: the static KV cache’s increased memory traffic more or less cancels out the reductions from fusing operations. By reducing per-weight memory footprint, AutoQuant reduces the memory traffic by 3.1$\times$ on top of SDPA and torch.compile. Applying LayerSkip on top of all system-level optimizations reduces the FLOPs count by 2.3$\times$ and the memory traffic by 2.2$\times$.
Figure 10: Operator Time Breakdown of Code Llama Rozière et al. (2024), Seamless Communication et al. (2023), Chameleon Team (2024a), HSTU Zhai et al. (2025) on H100 GPU.
4.5 Result Analysis over GPU Generations
In this section, we extend our analysis to NVIDIA H100 GPU NVIDIA (2023) to demonstrate how our insights and optimizations generalize across different hardware generations. H100 (Hopper) is the newer generation GPU beyond NVIDIA A100 (Ampere), introducing improvements in both computing capability and memory subsystem, achieving about 3$\times$ higher theoretical peak FLOPS and 1.5$\times$ higher HBM bandwidth compared to A100.
By examining the same workloads on both platforms, it brings insights on how the new generation of hardware could impact the performance characteristics in diverse aspects by showing how existing bottleneck is resolved with the architectural improvements and what new optimization opportunities come up. These cross-platform insights are crucial for both hardware architects designing future accelerators and system engineers optimizing software stack.
Figure 10 shows H100 GPU operator time breakdown compared to A100 (Figure 4). We observed two changes in performance characteristics. First, H100 demonstrates substantial improvements in computational efficiency, with Linear operations showing the most dramatic speedup of 6.82$\times$, while Attention operations achieve a 1.44$\times$ improvement resulting in 1.68$\times$ speedup for end-to-end baseline runtime for batch size 1. Second, significant acceleration in Linear operation has shifted the performance bottlenecks - models previously bounded by Linear operations now show Misc or Attention operations as their primary bottlenecks.
Figure 11: End-to-end inference time speedup with SDPA and SDPA+torch.compile for Llama and Chameleon on H100.
Figure 11 shows speedup with system-level optimizations. When all possible set of system-level optimization techniques are enabled for each workload, we get 2.21$\times$, 3.1$\times$, 1.5$\times$, 2.7$\times$ for Llama 34B, Chameleon 34B, Seamless (S-S) and HSTU for batch size 1 setting, respectively. LayerSkip on top of system-level optimizations gives the final speedup of 2.21$\times$, 4.13$\times$, 3.22$\times$, 4.53$\times$ speedup for T-T of Llama 34B & 7B, IT-T and I-T task of Chameleon 7B, respectively.
The reduced relative performance gains from optimizations on H100 compared to A100 can be attributed to enhanced baseline capabilities of H100. With architectural improvements including hardware-optimized attention mechanisms and higher memory bandwidth, it demonstrates diminishing returns in software optimization techniques as baseline hardware performance improves, despite the superior absolute performance of H100.
5 Key Lessons and Concluding Remarks